1000 recent changes in PartnerObserving Web retrieved at 01:48 (GMT)

SciPrep20110326
26 March 2011 LBC: continued testing with BP10 and checkout for science run ~1:30 John editing LBC header templates to match the latest master slave ones. CRPIX1,...
SciPrep20110324
Testing LBC with BP10 Plan is to run some science observations to test LBC guiding under BP10. Two fields: LQG, SN09. OBs in home/LBTO/Calib_OBs/ASTRONtests * ...
SciPrep20110309
2011 March 09 UT Technical Night. Plan To be done: * Check the Z4/Z11 relation now that the abovementioned changes have been made. * Run dofpia wi...
14Mar2011_Log
March 14, 2011 Observers: J. Bechtold, D. Thompson, F. Bian, L. Jiang 2:25pm. Lights off in the dome. Started series of 900s darks for Linhua's observations ton...
Get_data_from_archive
http://lbtarchive.as.arizona.edu/
ScienceMeeting2011
Meeting Web Site http://medusa.as.arizona.edu/lbto/LBT%20Conference/index.html Organization * SOC: Fan (Chair), Green/Wagner, one STC rep from each partner...
Check_Scripts
Check your scripts with Steven's perl script. Download the latest version of the script checker, e.g. lucichk3.2.pl (version 3.2) from the Partner Observing Page....
Copy_Lucifer_Data
Copy your data to your computer before you go so you don't have to fuss with the archive. The data files are on obs2 in /Repository/
Lucifer_Longslit
How to put a star in the Lucifer Long Slits Here's the idea: * Your script takes two direct images offset from each other and you difference them, to either ...
Imaging_flats
How to take imaging flats with the internal lamp * Be sure the dome is dark * Set up spectrograph in Lucifer ONE Control Window Select filter Gr...
ReadoutManagerGui
Lucifer ONE Readout Control Gui ReadoutManagerGui
Lucifer_Abort
To abort a Lucifer Script Type exit in the window you started the script in. or click on the big red ABORT! button on the ReadoutManagerGui to stop the curren...
Lucifer_Execute
To execute a Lucifer script 1. To start If your script is called myscript.txt Check the lower right corner of the Telescope Control window and select Wait for ...
FlushLucifer
Flush the Lucifer Detector If you went to bed and forgot to close up, or you observed a bright star, you may have to flush out the charge from the detector. * ...
Check_Scripts_On_Lucifer
Hopefully you checked your scripts at home with Steven's perl script, lucichk3.2.pl (or whatever the latest version is). You may also check your scripts on lucife...
CloseLucifer
At the end of the night * At a unix prompt on lucifer, type endofnight This moves FW1 = blind, FW2 = blind and M...
13Mar2011_Log
March 13, 2011 Observers: J. Bechtold, D. Thompson, F. Bian, L. Jiang Cloudy all afternoon 6:50pm. Sun has set. Completely overcast. 7:05pm Started two 15 min...
Cookbook_zJHK_Imaging
How to set up scripts for direct imaging with LUCIFER 1. Find a guide star for your target. 2. Create a script for acquiring your target and running a dither...
Lucifer_Start
Starting Up 1. Log on to obs2 as AZ, Click on Red LUCIFER icon to start guis if they aren't up from last night. password is written underneath the keyboard of ob...
Cookbook_ZJHK_Imaging
Scripts for zJHK Imaging
LuciferGuideStars
Find a good guide star for your target. Read the Lucifer Observation Preparation Suggestions page. Pick a guide star for your target Use Olga's perl script, lucig...
6Jan2011_Log
Thursday night, Jan. 6, 2011 Observers: Jill Bechtold, Dave Thompson Afternoon: Completely cloudy. Sunset: Fairly cloudy but will try Lucifer imaging. 5pm. Sw...
5Jan2011_Log
January 5, 2011 Observers: Jill Bechtold, Dave Thompson Completely overcast all afternoon and at sunset. Not looking good for tonight! LUCIFER's mask moving mec...
12Mar2011_Log
March 12, 2011 LUCIFER 1 Mask exchanger failed yesterday morning. Germans worked on it all day, but do not know what is wrong. We decided to do only imaging ton...
AZblog2011Mar
AZ March Run Nightly Observing Logs * Mar 10, 2011 * Mar 11, 2011 * Mar 12, 2011 * Mar 13, 2011 * Mar 14, 2011
11Mar2011_Log
Observers: Marie Rex, Wiphu Rujopakarn, Dave Thompson Too cloudy for twilight sky flats. 0.6" seeing. Taking test images with Br gamma filter detector looks sa...
10Mar2011_Log
Observers: Marie Rex, Wiphu Rujopakarn, Dave Thompson Clear, seeing .4 .5" 7:30 pm (2:30 UT) start 2 hours of H band science observations of MACS 0717 (Rex) Stil...
SciPrep20110307
LUCI prep/flexure testing: UT 2011 Mar 07 Observer: DThompson @LBT, OKuhn, ARakich in Tucson Telescope Operator: GBechetti Instrument Support: DOfficer Telesc...
SciPrep20110225
LUCI/LBC Technical Observing, prep for Partner Science: UT 2011 Feb 25 Observer: M.Pedani Telescope Operator: S.Allanson Instrument Support: T.Shih Telescope ...
SciPrep20110226
LUCI/LBC Technical Observing, prep for Partner Science: UT 2011 Feb 26 Observers: M.Pedani. A.Rakich,. O.Kuhn from downtown Telescope Operator: S.Allanson Inst...
SciencePrepNightLogs2010
Night Logs from 2010 28 30 Dec 2010 UT LUCIFER/LBC Science Prep (combined, mostly bad weather). 01 Nov 2010 UT LUCIFER/LBC Science Prep. 31 Oct 2010 UT LUCIFER/...
ObsInaf20110204
INAF LBC LUCIFER queue observing, 2011 February 03 04 Observer: K. Boutsia, R. Speziali Telescope Operator: S. Allanson Support Astronomer: O. Kuhn Instrume...
ObsInaf20110203
INAF LBC LUCIFER queue observing, 2011 February 02 03 Observer: K. Boutsia, R. Speziali Telescope Operator: S. Allanson Support Astronomer: O. Kuhn Instrume...
ObsInaf20110201
INAF LBC LUCIFER queue observing, 2011 January 31 February 01 Observer: K. Boutsia, R. Speziali Telescope Operator: S. Allanson Support Astronomer: D. Thomps...
ObsInaf20110202
INAF LBC LUCIFER queue observing, 2011 February 01 02 Observer: K. Boutsia, R. Speziali Telescope Operator: S. Allanson Support Astronomer: O. Kuhn Instrume...
ObsInaf20110131
INAF LBC LUCIFER queue observing, 2011 January 30 31 Observer: K. Boutsia, R. Speziali Telescope Operator: S. Allanson Support Astronomer: D. Thompson Instru...
SciPrep20110129
LBC/LUCI Technical Observing, prep for Partner Science: UT 2011 Jan 29 Observer: OKuhn (Tucson), DThompson (LBT) Telescope Operator: DHuerta Instrument Support...
SWTest20110128
LBT SW Testing and Range Balancing: UT 2011 Jan 28 Observer: DThompson Olga Kuhn and Andrew Rakich (Tucson) Telescope Operator: DHuerta Instrument Support: JMor...
10Jan2011_Log
January 10, 2011 OBSERVERS: Jill Bechtold, David Sand, Marco Pedani, Michael Rutkowski, Russell Ryan OPERATOR: David Gonzales Huerta Afternoon: Some high clouds....
AZLBCQuickReference
LBC Quick Reference Wiki Page AZ version, Jan 5, 2011 Start Up * Log into LBTO workstation (obs2) using your partner account (INAF, LBTB, OSURC, AZ). If yo...
9Jan2011_Log
January 9, 2011 OBSERVERS: Jill Bechtold, David Sand, Marco Pedani, Michael Rutkowski, Russell Ryan OPERATOR: David Gonzales Huerta Afternoon: Clouds. 3:43 pm....
AZblog2011Jan
AZ Jan Run Nightly Observing Logs * Jan 05, 2011 * Jan 06, 2011 * Jan 07, 2011 * Jan 08, 2011 * Jan 09, 2011 * Jan 10, 2011 * 2011011...
RunSummary_Jan2011
Run Summary: Arizona Partners, January 5 10, 2011 The mask exchange mechanism of Lucifer 1 failed just before the start of the AZ run, and the decision was made...
8Jan2011_Log
January 8, 2011 OBSERVERS: Jill Bechtold, David Sand, Marco Pedani, Michael Rutkowski, Russell Ryan OPERATOR: David Gonzales Huerta Our worries seem insignifica...
7Jan2011_Log
January 7, 2011 OBSERVERS: Jill Bechtold, Dave Thompson VERY CLEAR all afternoon. Preparing Lucifer. Changed Save path to /data/luci/20110108/ 3:48pm Took dark...
Lucifer_WFS
Collimate the Telescope January 2011: Operator will do this, you don't have to. 1. Find a Persson Standard near your first object. LOAD CATALOGUE LOAD ...
Lucifer_Pointing
Check Telescope Pointing by taking a Lucifer Science Image and running ptcorr Jan. 2011: At the beginning of the night, the Operator will go to a star near yo...
NoMask
Main.XiaohuiFan 08 Jan 2011
LookAtLuciferData
How to Look at your Lucifer Imaging during the night You can use the Skycat window to examine your data as it's being taken * Set up bias image in Skycat : C...
ScriptExamples
LUCIFER Script Examples This is a compilation of scripts that have been used to obtain LUCIFER data and illustrate its various modes. Jan. 2010: END_OSERVING_S...
ObserveTelluricStandard
How to Observe A Telluric Standard for a MOS mask We have written a perl script that allows placing the Telluric star in the leftmost and rightmost slits in an MO...
AZTargets2011Jan
Plan for Jan 2011 Run * Jan 5: Rex/Bechtold * Jan 6/7/8: Teske transit program Rex/Becthold * Jan 9/10: First half: Rex/Bechtold; second half: W...
AZObserverSchedule2011Jan
Observers: Jan 5 8: Bechtold, Rex, Teske Jan 9 10: Bechtold, Sand, Ryan Main.XiaohuiFan 02 Jan 2011
Pointing20101215
Generating an LBC Pointing Model from the LBC Red data taken on 20101215 The steps followed to generate a pointing model from these data taken on 20101215 are lis...
InafProjects2010a
First Column: C = 'completed' Ranking: A=4 5 B=3 4 C=2 3 LBC S PI Object Filter Rank Req. time tot. exec. time *UT da...
Projects2010bStatus
First Column: S(tatus) * C = 'completed' * P = 'partially executed' LBC *S *Object Filter Requested time Obs. time in specs Total Obs. t...
ObsInaf20101011
INAF LBC LUCIFER queue observing, 2010 October 10 11 Observer: V. Testa, K.Boutsia, M. Fumana Telescope Operator: D.Gonzales Huerta Support Astronomer: O.Kuhn I...
Log2010Nov4UT
skyflats begin at 12:30 (frames 01 0132, filters are brgam, h2, ks, h,j) standard acquisition script producing error. error is: error committing instrument setu...
Log2010Nov3UT
Night begins with Rex telluric and science (1:06) (1:30) Wiphu starts (3:50) and completes his program (5:30) seeing variable from 1.5 3" Back to Rex (5:45) see...
AZPlan2010Nov
Program Progress in Oct * Willmer, 6 hours of useful data, needs 3 4 more hours, target at 02h05d * McLinden, 3.3 hours of useful data, needs 40 min more, t...
Log2010Nov2UT
Observers: Marie, Emily, Wiphu, Greg Plan is to observe McLinden (~20 minutes), Rex (~4hrs), then Willmer for rest Observations begin with McLinden standard star...
AZblog2010Nov
AZ Nov Run Nightly Observing Logs (UT dates) * 2010 November 2 * 2010 November 3 * 2010 November 4 Main.EmilyMcLinden 02 Nov 2010
SciPrep20101101
LUCIFER/LBC Technical Observing, prep for Partner Science: UT 2010 Nov 01 Observer: DThompson, OKuhn in Tucson Telescope Operator: GBechetti, SAllanson Instrum...
SciPrep20101031
LUCIFER/LBC Technical Observing, prep for Partner Science: UT 2010 Oct 31 Observer: DThompson, OKuhn in Tucson through 2:30am Telescope Operator: GBechetti, SAl...
Log2010Oct14UT
Pointing problems to begin the night McLinden acquisition and telluric additional integrations needed to find ref stars, seeing bad 1.5" Begin shortened McLi...
AZblog2010Oct
AZ Oct Run Nightly Observing Logs (UT dates) * 2010 October 12 * 2010 October 13 * 2010 October 14 Observing Log: * 2010 October 12 Main.EmilyMc...
AZPlan2010Oct
AZ Oct Run Oct 11, 12, 13, LUCIFER, Observers: Emily McLinden, Marie Rex, Christopher Willmer Oct 11: First half: McLinden/Rieke, Second half: Willmer Oct 12: ...
Log2010Oct13UT
Observers: Emily McLinden, Marie Rex, Christopher Willmer, Olga Kuhn, Steve Allanson, Gino Bechetti Mirror temperature about 4 degrees C warmer than yesterday. 12...
Log2010Oct12UT
Summary : Clear night, Seeing ~ 0.5". A significant portion of time was lost because of several mishaps collimation failing, not finding objects, script errors...
r4 - 12 Oct 2010 - 10:59 by ChristopherWillmer
AzObservingLog2010Oct12UT
* *observers mclinden kuhn rex willmer * *Date UT 20101012 file numbers object Camera * *Grating Filter1 Filter2 DIT ...
r2 - 12 Oct 2010 - 07:17 by ChristopherWillmer
2010Oct12UT
Observers : McLinden? , Rex, Willmer, Kuhn, David Lucifer MOS SkyFlats? by accident some of the sky flats were given an incorrect date. Flats 20101011_Flat_00...
NEW - 12 Oct 2010 - 03:02 by ChristopherWillmer
ObsInaf20101010
INAF LBC LUCIFER queue observing, 2010 October 09 10 Observer: V. Testa, K.Boutsia, M. Fumana Telescope Operator: D.Gonzales Huerta Support Astronomer: O.Kuhn I...
ObsInaf20101009
INAF LBC LUCIFER queue observing, 2010 October 08 09 Observer: V. Testa, K.Boutsia, M. Fumana Telescope Operator: D.Gonzales Huerta Support Astronomer: O.Kuhn I...
ObsInaf20101008
INAF LBC LUCIFER queue observing, 2010 October 07 08 Observer: V. Testa, K.Boutsia, M. Fumana Telescope Operator: D.Gonzales Huerta Support Astronomer: O.Kuhn I...
ObsInaf20101007
INAF LBC LUCIFER queue observing, 2010 October 06 07 Observer: V. Testa, K.Boutsia, M. Fumana Telescope Operator: D.Gonzales Huerta Support Astronomer: O. Kuhn ...
LMSScriptGen
Lucifer Acquisition and Spectroscopic Script Generator (lmsscriptgen) * Steve Allanson, LBTO (05 October 2010) Thank you for volunteering to beta test this scr...
SciPrep20100929
LUCIFER Technical Observing, prep for Partner Science: UT 2010 Sep 29 Observer: DThompson Telescope Operator: SAllanson, GBechetti Instrument Support: TShih Te...
SciPrep20100919
19 September 2010 Afternoon: * Assessed focal plane tilt. I suspect a small ~50 micron, tilt in Blue, such that chip 1 is closer to M1 than chip 3. But need t...
SciPrep20100918
18 September 2010 2nd half LBC restart * Doug Geno plugged in the cable that was brought to Tucson and repaired by Dave Baxter. All four chips on LBC red are...
SciPrep20100917
17 September 2010 2nd half * LBC Chip 2 still dead. See lbcr.20100917.071852.fits. It is not responding to starlight, but has only bias counts plus crosstalk f...
MakingSlitmasks
Masks are due by noon in Tucson on the following dates for 2010B: * 7 September Arizona (8), INAF (6), OSU/RC (9) * 10 October Arizona (11), LBTB (11); (the...
SciPrep20100914
LFBG/LUCIFER Restart after shutdown: UT 2010 Sep 14 Observer: Telescope Operator: Instrument Support: Telescope Support: SW support: Summary Preparation...
SciPrep20100606
LUCIFER Technical Observing, prep for Partner Science: UT 2010 Jun 06 Observer: DThompson Telescope Operator: SAllanson Instrument Support: JMorris Telescop...
SciPrep20100702
LUCIFER Training Preparations for Partner Science: UT 2010 July 02 Observer: DThompson, KRueff, PHibon Telescope Operator: SAllanson Instrument Support: TShi...
SciPrep20100701
LUCIFER Training Preparations for Partner Science: UT 2010 July 01 Observer: DThompson, KRueff, PHibon Telescope Operator: SAllanson Instrument Support: TShi...
Projects2010aStatus
First Column: S(tatus) * C = 'completed' * P = 'partially executed' LBC *S *Object Filter Requested time Obs. time in specs Total Obs. t...
r11 - 22 Jun 2010 - 09:32 by VincenzoTest
ObsInaf20100617
INAF LBC LUCIFER queue observing, 2010 June16 17 Observer: V. Testa, K.Boutsia Telescope Operator: S.Allanson Support Astronomer: O. Kuhn Instrument: LUCIFER1 ...
r6 - 22 Jun 2010 - 09:16 by VincenzoTest
ObsInaf20100618
INAF LBC LUCIFER queue observing, 2010 June17 18 Observer: V. Testa, K.Boutsia Telescope Operator: S. Allanson Support Astronomer: O. Kuhn Instrument: LUCIFER1 ...
r5 - 22 Jun 2010 - 09:10 by VincenzoTest
ObsInaf20100616
INAF LBC LUCIFER queue observing, 2010 June15 16 Observer: V. Testa, K.Boutsia Telescope Operator: S.Allanson Support Astronomer: O. Kuhn Instrument: LUCIFER1 ...
r4 - 22 Jun 2010 - 08:53 by VincenzoTest
ObsInaf20100613
INAF LBC LUCIFER queue observing, 2010 June 12 13 Observer: V. Testa, K.Boutsia Telescope Operator: D.Gonzales Huerta Support Astronomer: M. Pedani Instrument: ...
r8 - 22 Jun 2010 - 08:00 by VincenzoTest
PostSciDebrief2010June
Issues which arose during the 7 18 June 2010 Science Run LBC: FPIA using incorrect configuration file Observers had a copy of the file lbcfpia.cfg in their ho...
ObsInaf20100615
INAF LBC LUCIFER queue observing, 2010 June14 15 Observer: V. Testa, K.Boutsia Telescope Operator: D.Gonzales Huerta Support Astronomer: O. Kuhn Instrument: LUC...
r6 - 16 Jun 2010 - 07:23 by VincenzoTest
ObsInaf20100614
INAF LBC LUCIFER queue observing, 2010 June 13 14 Observer: V. Testa, K.Boutsia Telescope Operator: D.Gonzales Huerta Support Astronomer: O. Kuhn Instrument: LU...
r6 - 15 Jun 2010 - 10:07 by VincenzoTest
DuplicateLuciMaskChecks
How to check the duplicate pairs of masks in LUCIFER, June 2010 We are assessing the quality of cuts produced by the LBTO laser milling machine. This is done usi...
SciPrep20100605
LBC Preparations for Partner Science: UT 2010 Jun 05 Observers: OKuhn, DThompson, MPedani (LBTO) Telescope Operator: SAllanson Instrument Support: JMorris Sum...
20100508logUT
LBC Tech Chip IA Commissioning 08 May 2010 UT Night 4 Observers: M.Pedani (LBTO) Telescope Operator: SAllanson (LBTO), JHill (LBTO) Telescope Support: JUrban(...
TechNightLogs
LBC Technical Nights 2010 02 January UT (January Tech Night 1) Collimation and pointing data 27 January UT LBC Blue Active Optics Tests in second half 30 Ja...
20100507logUT
LBC Tech Chip IA Commissioning 07 May 2010 UT Night 3 Observers: M.Pedani (LBTO) Telescope Operator: SAllanson (LBTO), JHill (LBTO) Telescope Support: JUrban(...
20100505logUT
LBC Tech Chip IA Commissioning 05 May 2010 UT Night 1 Observers: M.Pedani (LBTO) Telescope Operator: SAllanson (LBTO), JHill (LBTO) Telescope Support: DAshby...
20100506logUT
LBC Tech Chip IA Commissioning 06 May 2010 UT Night 2 Observers: M.Pedani (LBTO) Telescope Operator: SAllansen (LBTO), JJHill (LBTO) Telescope Support: Midki...
SciPrep20100501
LUCIFER Training Preparations for Partner Science: UT 2010 May 01 Observer: DThompson (LBTO) Telescope Operator: WWack Instrument Support: TShih Telescope Su...
SciPrep20100504
LUCIFER Training Preparations for Partner Science: UT 2010 May 04 Observer: DThompson (LBTO) Telescope Operator: WWack (Tucson) Instrument Support: TShih Tel...
SciPrep20100502
LUCIFER Training Preparations for Partner Science: UT 2010 May 02 Observer: DThompson (LBTO) Telescope Operator: WWack (Tucson) Instrument Support: TShih Tel...
ObsInaf20100411
INAF LBC LUCIFER queue observing, 2010 April10 11 Observer: A. Fontana, K.Boutsia, R. Speziali Telescope Operator: D.Gonzales Huerta Support Astronomer: D. Thomp...
Log100418
Summary: Too much cloud to do anything scientific (in fact, the dome was closed from 4:50 to 8:20). The night was however useful to practice various LUCIFER obse...
AZblog2010Apr
AZ Apr Run Nightly Observing Logs (UT dates) * April 15, 2010 * April 16, 2010 * April 17, 2010 * April 18, 2010 Main.EiichiEgami 18 Apr 2010
Log100417
Summary: Heavy cloud cover at the beginning, but cleared up enough to start observing at 3:40 (but had some clouds most of the time). Dome again closed for 2.5 ...
Log100416
Summary: Sky was clear initially, but clouds started to move in around 5:30 UT. Seeing was sub arcsecond at the beginning, but quickly deteriorated to 1.5" 2". ...
Log100415
Summary: Some clouds at the beginning but cleared up later. Seeing was mediocre initially (~1") but improved to ~0.8". We have executed 2 OBs for E. Olszewski's...
AZPlan2010Apr
AZ April run Apr 14, Apr 15 , LBC, Observers: Jiang, Egami Apr 16, Apr 17, LUCIFER, Observers: Egami, Rieke team On LBC nights: Before 2300 local time: Ol...
ObsInaf20100414
INAF LBC LUCIFER queue observing, 2010 April13 14 Observer: K.Boutsia, R. Speziali Telescope Operator: D.Gonzales Huerta Support Astronomer: O. Kuhn Instrument...
r7 - 14 Apr 2010 - 21:01 by VincenzoTest
ObsInaf20100409
INAF LBC LUCIFER queue observing, 2010 April08 09 Observer: A. Fontana, K.Boutsia, R. Speziali Telescope Operator: D.Gonzales Huerta Support Astronomer: D. Thomp...
r11 - 14 Apr 2010 - 13:02 by VincenzoTest
ObsInaf20100410
INAF LBC LUCIFER queue observing, 2010 April09 10 Observer: A. Fontana, K.Boutsia, R. Speziali Telescope Operator: D.Gonzales Huerta Support Astronomer: D. Thomp...
r12 - 14 Apr 2010 - 13:01 by VincenzoTest
ObsInaf20100412
INAF LBC LUCIFER queue observing, 2010 April11 12 Observer: R. Speziali, K.Boutsia Telescope Operator: D.Gonzales Huerta Support Astronomer: D. Thompson Instrum...
r6 - 14 Apr 2010 - 12:58 by VincenzoTest
ObsInaf20100413
INAF LBC LUCIFER queue observing, 2010 April12 13 Observer: K.Boutsia, R. Speziali Telescope Operator: D.Gonzales Huerta Support Astronomer: O.Kuhn Instrument: ...
r3 - 14 Apr 2010 - 10:41 by VincenzoTest
AZObserverNote2010Apr
Notes to AZ Observers About the Queue * First two nights on LBC, last two nights on LUCIFER * In marginal conditions, use Milne program as backup * In c...
AZTargets2010Apr
PI Instrument Object RA DEC Mode Filter Grating Exp Time Status Jiang LBC IRAC Dark Field 17:40:00 69...
20100408logUT
LBC Technical Observations 8 April 2010 UT 07:30 Starting LBC observing. LUCIFER technical observing was done during the first half 07:32 Focus/Collimate on ACT02...
ObsInaf20100218
INAF LBC LUCIFER queue observing, 2010 February17 18 Observer: V. Testa, K.Boutsia, F. Pedichini Telescope Operator: D.Gonzales Huerta Support Astronomer: D. Tho...
r5 - 09 Apr 2010 - 06:27 by VincenzoTest
SciPrep20100408
LBC LUCIFER Preparations for Partner Science: UT 2010 Apr 08 Observer: OKuhn, MPedani (Tucson), DThompson (LBTO) Telescope Operator: David Gonzalez Huerta Instr...
SciPrep20100407
LBC LUCIFER Preparations for Partner Science: UT 2010 Apr 07 Observer: OKuhn, MPedani (Tucson), DThompson (LBTO) Telescope Operator: David Gonzalez Huerta Instr...
20100406logUT
7 April 2010 UT Technical Night 1 David Gonzalez Huerta, Dave Thompson, John Morris (LBTO) Marco Pedani, Olga Kuhn, Andrew Rakich (Tucson) Start with LBC 02:...
MoreExampleScripts
Attached is a tarball with example scripts for observing a target, including JHK imaging, acquiring an object in a long slit, observing a spectrum, and observing ...
LBCCollimationModels
LBC Collimation Models The table below will be used to store relevant facts about the recent collimation models for LBC SX and DX. * Column 1: the date on wh...
20100127UT
27 January 2010 LBC Blue Active Optics Tests during second half Operator/Observer: JHill, assisted by DMiller Summary The first half of this night is covered i...
ObsInaf20100219
INAF LBC LUCIFER queue observing, 2010 February18 19 Observer: V. Testa, K.Boutsia, F. Pedichini Telescope Operator: D.Gonzales Huerta Support Astronomer: D. Tho...
r7 - 03 Mar 2010 - 10:31 by VincenzoTest
20100217UT
Parallel Testing of LBC Red and DX Cell during LUCIFER observing 17 February 2010 UT Observer: JHill (Tucson) Summary JMH completed the first phase of the on sk...
ObsInaf20100216
INAF LBC LUCIFER queue observing, 2010 February15 16 Observer: V. Testa, K.Boutsia, F. Pedichini Telescope Operator: D.Gonzales Huertak Support Astronomer: D. Th...
r6 - 19 Feb 2010 - 06:49 by VincenzoTest
ObsInaf20100217
INAF LBC LUCIFER queue observing, 2010 February16 17 Observer: V. Testa, K.Boutsia, F. Pedichini Telescope Operator: D.Gonzales Huerta Support Astronomer: D. Tho...
r5 - 19 Feb 2010 - 06:41 by VincenzoTest
ObsInaf20100215
INAF LBC LUCIFER queue observing, 2010 February14 15 Observer: V. Testa, K.Boutsia, F. Pedichini Telescope Operator: WWack Support Astronomer: D. Thompson Instr...
r7 - 16 Feb 2010 - 02:36 by VincenzoTest
2010Feb15-16
Main.VincenzoTest 16 Feb 2010
NEW - 16 Feb 2010 - 01:21 by VincenzoTest
ObsInaf20100214
INAF LBC LUCIFER queue observing, 2010 February13 14 Observer: V. Testa, K.Boutsia Telescope Operator: WWack Support Astronomer: D. THompson Instrument: LBC S...
Log100212
01:00 The sky looks nice, and we are open. 01:30 check pointing. met a problem when running ptcorr, 'subsystem not running failed', ptcorr can not send to positio...
AZTargets2010Jan
PI Priority Object RA DEC Mode Filter Grating Exp Time Status Fan A J2227 0605 22:27...
Log100211
Almost clear, only some clouds around the horizon. 00:30 Dome is open to cool down the telescope. 00:50 The humidity is going up now is 60%. 01:10 The humidity ...
Log100210
00:30 It is snowing. no open 09:00 It is getting clear, but the humidity is 100%. 13:00 No open in the whole night due to the snow and high humidity. Main.Fuyan...
Log100209
12:57 There are thick clouds around us, and thin clouds on the top of the sky. The Humidity is 100%. Main.FuyanBian 10 Feb 2010
AZblog2010Feb
AZ Feb Run Nightly Observing Logs (UT dates) * Feb 09, 2010 * Feb 10, 2010 * Feb 11, 2010 * Feb 12, 2010 Main.FuyanBian 10 Feb 2010
AZPlan2010Feb
AZ Feb run (UT Time) Feb 10, Feb 11: LUCIFER, Observers: Bian, McGreer, Hibon Feb 12, Feb 13: LBC, Observers: Bian, Just LUCIFER Time (weather looks pretty hopel...
20100131UT
31 January 2010 Technical Night 2 continued LBC Active Optics testing LUCI prep (20100131) Wayne Wack, Dave Thompson at LBT Andrew Rakich, Olga Kuhn at Stewar...
20100130UT
30 January 2010: Active Optics tests 05:10 Open. Cloudy skies, but humidity dropped below threshold 05:13 Slew to ACT0370, pointing star. Some bumps with observin...
ObsInaf20100118
INAF LBC LUCIFER queue observing, 2010 January 17 18 Observer: V. Testa, F. Mannucci Telescope Operator: WWack Instrument: LBC Summary * Time lost: ~8h lost fo...
ObsInaf20100119
INAF LBC LUCIFER queue observing, 2010 January 17 18 Observer: V. Testa, Telescope Operator: WWack Instrument: LBC Summary * Time lost: 1:15 for clouds at the...
ObsInaf20100117
INAF LBC LUCIFER queue observing, 2010 January 16 17 Observer: V. Testa, F. Mannucci Telescope Operator: WWack Instrument: LBC Summary * Thin and thick clouds ...
LogBook20100116
automatically created by IRtools/obssum File Object * RA* Dec Filt Co x Texp * Airm* * UT_obs* 1 dark 15:39:00...
ObsInaf20100116
INAF LBC LUCIFER queue observing, 2010 January 15 16 Observer: V. Testa, F. Mannucci, D. Thompson Telescope Operator: WWack Instrument: LUCIFER Data filename roo...
LogBook20100115
automatically generated by IRtools/obssum File Object * RA* Dec Filt/Gra * Co x Texp* Airm UT_obs 1 undefined 8:0...
ObsInaf20100115
INAF LBC LUCIFER queue observing, 2010 January 14 15 Observer: V. Testa, F. Mannucci, D. Thompson Telescope Operator: WWack Instrument: LUCIFER Data filename roo...
Log100109
Summary We have around 2 hours science data, most of them are useless due to the cloud. We also have 2 hours fun data, e.g. abell cluster, globular cluster, inter...
FeII
Main.FuyanBian 09 Jan 2010
Log100108
Summary It is very cloudy tonight, we lost 5 hours due to the bad weather, 2.5 hours due to the telescope, which is the pointing, we can not find the guide star f...
2009Sep11UT
LBC on 11 September 2009 UT Nautical Twilight at 02:26 UT We are closed due to threatening clouds. The LBC team tried to put some new software on CMU without succ...
Log100107
Summary It is cloudy tonight. Totally, 2~hours scientific data taken. J0203 0012 H K spectra 40min SN 2008hy J H imaging 34.5min 50.0min J104226 H K spectra...
AZLuciExperience
This page is for notes on user experiences with Lucifer. * Selection of guide stars is a consistent problem. The sky area available for guide stars is quite s...
Log100106
The sky is clear. Summary There are total ~2.8 hours scientific exposure. 32 min J0900 2234 J H K image 25 min J0900 2234 H K spectra (bad seeing) 30 min 3C220 ...
Log100105
All times UTC. Flats 12:14 Started continuous images in manual mode with Br_gam and 1.98s integration. Level ~5000. 12:17 Set to Ks and 1.98s. Level ~30,000. ...
AZplan2010Jan
AZ Jan Run Nightly Observing Plan (UT dates) * Jan 05, 2010 * Jan 06, 2010 * Jan 07, 2010 * Jan 08, 2010 * Jan 09, 2010 Main.FuyanBian 03...
AZblog2010Jan
AZ Jan Run Nightly Observing Logs (UT dates) * Jan 05, 2010 * Jan 06, 2010 * Jan 07, 2010 * Jan 08, 2010 * Jan 09, 2010 Main.XiaohuiFan 0...
AZObserverNote
Notes to AZ Observers About the Queue * Do priority A objects first * J2227 is probably too low, so for the program, please observe one or two of the backu...
AZObserverSchedule2010Jan
AZ Jan Block Observer Schedule Date Observer Jan 4 8 Fuyan Bian Jan 4 6 Ian McGreer Jan 6 8 Jill Bechtold Jan 7 8 Emily McLi...
LUCSDTblog
LUCIFER SDT Nightly Observing Logs (UT dates) * Engineering 20091205.log: Log from Dec 5, 2009 * Dec 6, 2009 20091206.log: Log from Dec 6, 2009 * D...
Log091217
17 December 2009. Last night of SDT LUCIFER observing. At sunset there is light cirrus. This will affect our flats, but we will take some at J and K anyway. Note,...
LUCSDTPartner
Partner Hours Completed During SDT Partner Hours Comments AZ 2.7 INAF 2.9 LBTB 1.9 OSURC 2.0 Total 9.4 includes 0....
LUCSDTProgress
Targets Completed PI Partner Priority Object Mode Filter Grating Exp Time Comments Date Germeroth LBTB A SDSSJ0222 0003 Imaging ...
ApprovedTargetsA
Approved LUCIFER SDT Program : Priority A PI Partner Priority Object RA DEC Mode Filter Grating Exp Time...
Log091216
16 December 2009 (UT=LST 1:40) Queue Observers Chick W.,Olga K., Andrea Q., Roberto S. Some high cirrus. But we are able to open after being shut. FLATS 16...
Log091215
Telescope closed due to ice blocks on the roof. The new obs team performed some practice simulating observations and testing scripts with the dome closed.
Log091211
Main.GiovanniCresci 11 Dec 2009 Observers: Manolis, Giovanni, Dave, Nancy, Walter, Marco, Olga 00:20 UT sky partly cloudy at sunset, with clouds moving towards...
Log091214
UT0100: We are closed due to 100% humidity and threatening clouds. UT0500: Snow has started to fall. Paul Martini 14 Dec 2009
Log091213
UT 03:07 Sky is almost clear, but 100% humidity. Not open due to high humidity. Closed all night due to high humidity. Fuyan Bian/Paul Martini 13 Dec 2009
Log091212
Sunset: The humidity is 85% and the trend is not favorable, we are therefore are not opening. UT 02:00 The humidity is 100%, although the sky is fairly clear. S...
BadWeatherProgram
The following objects are relatively bright targets among approved programs. Observers should consult proposal and observing files to decide if they need good ima...
LUCtips
Preparing LUCIFER Scripts While reviewing some of the LUCIFER SDT scripts, we (Ian Jochen) found a couple of common mistakes: Absolute offsets are based on a ...
Log091210
Main.GiovanniCresci 10 Dec 2009 Telescope closed due to ice blocks on the roof.
Log091209
Telescope closed due to snow accumulation on roof. No observing. Main.XiaohuiFan 09 Dec 2009
Log091208
LUCIFER Science Demonstration Time: UT 2009 December 08 Observer: D. Thompson (@LBT), I. McGreer, J.Heidt (both went back to Tucson after a Blizzard warning for ...
Plan091208
No specific plan made, foggy/cloudy at afternoon, heavy snow predicted during the night (which happened indeed) Main.JochenHeidt 08 Dec 2009
Plan091207
As in the last night, mostly cloudy at beginning of night, strong winds close to 20m/s not clear if and when we can open = No detailed observing plan made. If it...
Log091207
LUCIFER Science Demonstration Time: UT 2009 December 07 Observer: D. Thompson, J. Heidt, I. McGreer (all@LBT) Telescope Operator: A. Ceranski Support: N. Ageorg...
Log091206
LUCIFER Science Demonstration Time: UT 2009 December 06 Observer: D. Thompson, J. Heidt, I. McGreer (all@LBT) Telescope Operator: A. Ceranski Support: N. Ageorg...
Plan091206
Cloudy at beginning of night, not clear if and when we can open = No detailed observing plan made. If it clears up and observng conditions are good, we'll start ...
LUCSDTplan
LUCIFER SDT Nightly Observing Observing Plans (UT dates) * Dec 6, 2009 * Dec 7, 2009 * Dec 8, 2009 * Dec 9, 2009 * Dec 10, 2009 * Dec 11, 2009 ...
ApprovedTargetsB
Approved LUCIFER SDT Program : Priority B PI Partner Priority Object RA DEC Mode Filter Grating Exp Tim...
LUCSDTObserverJob
SDT Observers' Tasks A. Queue Planning * In the afternoon, draft tonight's observing plan, including back up targets, and calibration plan based on conditions...
LUCObserverSchedule
LUCIFER SDT Observer Schedule Date Observer#1 Observer#2 Supporting Astronomer LUCIFER Team Dec 5 7 Ian McGreer (AZ) Jochen Heidt (LBTB) Thom...
2009Oct24UT
24 October 2009 LBC Tech Night 2 We took a lot of data last night and much still needs to be analyzed, but we learned some things right away. 1) Red Z22 zero poi...
2009Oct23UTl
2009 October 23 LBC Tech Night (active optics work see PLAN) Times are UT: 00:00 Took a series of 25 biases with Blue/Red these appear in newdata and in Repos...
2009Oct23UT
Archive work on 22 October 2009 UT The LBC data flow was changed (Riccardo Dave) so that the data go directly to /newdata which is mounted on the observer works...
ObservingLogs
OBSERVING LOGS for LBC SDT 2007 If you use WYSIWYG and "Formated" instead of "Normal" your logs will look ... well... "normal" jill January 12 24 SDT Run ...
2009Oct14UT
Second LBC Tech Night Data on:/Repository/20091014_Tech Observers: Testa, V, Pedichini F. for LBC team, Pedani, M. for LBTO Cirrus clouds T_air = 5.0 RH~65% Wi...
AGWHomePage
AWG Acquisition, Guiding and Wavefront sensing units * http://www.aip.de./groups/lbt/ Main.JorgeEsguerra 15 Oct 2009
NEW - 15 Oct 2009 - 05:40 by JorgeEsguerra
MODSHomePage
MODS Multi Object Double Spectrograph * http://www.astronomy.ohio state.edu/MODS/ Main.JorgeEsguerra 15 Oct 2009
NEW - 15 Oct 2009 - 05:37 by JorgeEsguerra
PEPSIHomePage
PEPSI Potsdam Echelle Polarimetric and Spectroscopic Instrument * http://www.aip.de/pepsi/ Main.JorgeEsguerra 15 Oct 2009
NEW - 15 Oct 2009 - 05:35 by JorgeEsguerra
LBTIHomePage
Large Binocular Telescope Interferometer * http://lbti.as.arizona.edu/LBTI Main/Project.html Main.JorgeEsguerra 15 Oct 2009
NEW - 15 Oct 2009 - 05:31 by JorgeEsguerra
LINCHomePage
LINC Nirvana * http://www.mpia hd.mpg.de/LINC/ Main.JorgeEsguerra 15 Oct 2009
NEW - 15 Oct 2009 - 05:30 by JorgeEsguerra
2009Oct13UT
First LBC tech Night. LBC team present. Nautical Twilights: 01:42 UT 12:28 UT 01:42 Closed due to strong winds and high RH 04:05 Weather still bad with RH~88% ...
SDTMtnTechSupport
Technical and Software Support for SDT Observations During the January SDT run, the following persons will be providing 24/7 telescope technical support. The des...
r7 - 09 Oct 2009 - 22:28 by NormCushing
2009Sep14UT
LBC on 14 September 2009 UT Clouds. 05:10 We open dome (still some thunderstorms south of Wilcox) T_air=9.4 RH=80% T_mirr=11.8 Right Shutter alarm! Left Opened...
2009Sep13UT
At 21:32 Sept 12 UT we start cameras to check with 10 Biases LBCB Mean r.m.s = 5.21 ADU LBCR Mean r.m.s = 5.46 ADU Red channel a little more noisy with some pick ...
2009Sep12UT
No Observing tonight. A hail storm in the afternoon produce half an inch of hail accumulated on the dome. We had serious water leaks that affected the LBC Tree Ho...
2009Sep05UT
05 September 2009 LBC Checkout Observers: O. Kuhn, J. Hill (from Tucson) Telescope Operator: D. Gonzalez Huerta Summary Lost to weather. Preparations * TCS B...
2009Sep04UT
4 September 2009 LBC Checkout Observer: O. Kuhn Telescope Operator: D. Gonzalez Huerta Test L2 position Check Focus (z SLOAN) Install and test new mcspu Runni...
2009Sep02UT
02 September 2009 (First night of LBC post shutdown checkout) 02:17 MST LBCs up and running. Note that during the afternoon HK blue readings had dropped out. So I...
SiteInfo
Sky Brightness * Sky Brightness and Seeing Measurements from data obtained at the VATT were published in Vatican Advanced Technology Telescope UBVR Sky Surfac...
18Mar2007
PartnerObserving Log for 2007 Mar 18/19 Observers: Rick Pogge, Eichii Egami Support Scientist: none TO: Aaron Ceranski Lots of high clouds around during the after...
r9 - 21 Jul 2009 - 22:55 by NormCushing
19Apr2007
PartnerObserving Log for 2007 Apr 19/20 Observers: A. Grazian, R. Gredel. Night Assistant: D. Thompson, J. Hill Telescope Operator: A. Ceranski 17:59 Turn on Hous...
r6 - 21 Jul 2009 - 22:55 by NormCushing
19Feb2007
PartnerObserving Log for 2007 Feb 19/20 Observer1: Nicolas Bouche, Peter Buschkamp Observer2: Paul Smith Inst. Support: Olga Kuhn TO: Aaron Ceranski Closed for S...
r3 - 21 Jul 2009 - 22:55 by NormCushing
19Jun2007
PartnerObserving Log for 2007 June 19 20 Observer: D. Thompson LBC Support: A. DiPaola, R. Speziali, R. Bertram Telescope Operator: D. Gonzalez 02:56 Dome open, p...
r3 - 21 Jul 2009 - 22:55 by NormCushing
19Mar2007
PartnerObserving Log for 2007 Mar 19/20 Observers: Rick Pogge, Eichii Egami Support Scientist: none TO: Aaron Ceranski Mostly clear during the afternoon, with hig...
r5 - 21 Jul 2009 - 22:55 by NormCushing
15Sep2007
PartnerObserving Log for 2007 Sep 15 16 Observers: J. Hill, O. Kuhn Telescope Operator: A. Ceranski LBTO Family night, but cloudy until midnight. Ray and Tim and ...
r4 - 21 Jul 2009 - 22:55 by NormCushing
16Apr2007
PartnerObserving Log for 2007 April 16/17 Observers: J. Harris, D. Sand Instrument Support: Telescope Operator: Instrument: LBC Blue $ 19:00: Camera shutter ...
r5 - 21 Jul 2009 - 22:55 by NormCushing
16Feb2007
PartnerObserving Log for 2007 Feb 16/17 Observer: Tom Herbst, Giacomo Beccari Inst. Support: Dave Thompson TOs: David Gonzalez, Aaron Ceranski 18:05 Mostly cle...
r8 - 21 Jul 2009 - 22:55 by NormCushing
16Jun2007
PartnerObserving Log for 2007 June 16/17 Observers: V. Burwitz, J. Hill Telescope Operator: A. Ceranski Instrument: LBC Blue Weather evening: thin clouds ...
r6 - 21 Jul 2009 - 22:55 by NormCushing
16Mar2007
PartnerObserving Log for 2007 Mar 16/17 Observers: Rick Pogge, Peter Garnavich Support: John Hill on site, Olga Kuhn on call TO: Aaron Ceranski Afternoon weather...
r9 - 21 Jul 2009 - 22:55 by NormCushing
16Oct2007
PartnerObserving Log for 2007 Oct 16 17 Observers: D. Sand and M. Coleman Telescope Operator: A. Ceranski Main.JohnHill 17 Oct 2007 Sand and Coleman here. ...
r5 - 21 Jul 2009 - 22:55 by NormCushing
17Apr2007
PartnerObserving Log for 2007 April 17/18 Observers: J. Harris, D. Sand Instrument Support: D. Thompson, R. Bertram Telescope Operator: Instrument: LBC Blue $...
r6 - 21 Jul 2009 - 22:55 by NormCushing
17Jun2007
PartnerObserving Log for 2007 June 17/18 Observers: V. Burwitz, J. Hill Telescope Operator: A. Ceranski Instrument: LBC Blue Weather evening: Clear Weat...
r6 - 21 Jul 2009 - 22:55 by NormCushing
17Mar2007
PartnerObserving Log for 2007 Mar 17/18 Observers: Rick Pogge, Eichii Egami Support: John Hill on site, Olga Kuhn on call TO: Aaron Ceranski Afternoon weather is ...
r8 - 21 Jul 2009 - 22:55 by NormCushing
18Apr2007
PartnerObserving Log for 2007 Apr 18/19 Observers: A. Grazian, R. Gredel. Night Assistant: D. Thompson, J. Hill Telescope Operator: A. Ceranski 18:01 Turn on Hous...
r5 - 21 Jul 2009 - 22:55 by NormCushing
18Feb2007
PartnerObserving Log for 2007 Feb 18/19 Observer1, Nicolas Bouche, Peter Buschkamp Observer2, Paul Smith Inst. Support, Dave Thompson TO; Aaron Ceranski UT 03:0...
r7 - 21 Jul 2009 - 22:55 by NormCushing
18Jun2007
PartnerObserving Log for 2007 June 18 19 Observer: D. Thompson Telescope Operator: D. Gonzalez 02:42 Crew still cleaning up in the dome. Dewar warmed sign...
r3 - 21 Jul 2009 - 22:55 by NormCushing
ObservingPreparation
Preparing for your PartnerObserving Run Planning your PartnerObserving Program * Generation of PartnerObserving Blocks (LBC only) * MakingSlitmasks for LUCI...
r2 - 21 Jul 2009 - 22:55 by NormCushing
October21UT
21 October 2008 PartnerObserving Log Copied new version which Stefano sent of track_first.exe and track_next.exe to /images on CMU. Moved existing versions to tr...
r6 - 21 Jul 2009 - 22:55 by NormCushing
SDTObsBlockUpload
SDT PartnerObserving Block Uploads Please email final OBs to okuhn,dthompson at as.arizona.edu so we can coordinate getting them up to the telescope in advance...
r3 - 21 Jul 2009 - 22:55 by NormCushing
September24UT
Afternoon * State of things: Spare controller which was swapped in earlier last night remains in place (ID #1). But we got the same error message about a bad ...
r8 - 21 Jul 2009 - 22:55 by NormCushing
SkyFlatCalib
Procedures for taking Twilight Sky Flats 1.After the telescope operator opens the dome (just after sunset) and the telescope and instrument are ready, ask him...
r3 - 21 Jul 2009 - 22:55 by NormCushing
SkyInLBCETC
Main.jill 12 Dec 2006 * NOTES on SKY VALUES IN ETC 12 Dec 2006* Jill Bechtold 11 December 2006 Revised, Dec. 22, 2006 SKY BRIGHTNESS OBSERVED V. THE LBT ETC Sum...
r3 - 21 Jul 2009 - 22:55 by NormCushing
12Mar2007
PartnerObserving Log for 2007 Mar 12/13 Observers: Rick Pogge Inst. Support: John Hill TO: Ed Lundin UTC Mar 12 2100: Fired up the system, cryostat temperature ...
r6 - 21 Jul 2009 - 22:55 by NormCushing
12May2007
PartnerObserving Log for 2007 May 12/13 MST Technical Observers: J. Hill, (A. DiPaola, R. Speziali) Telescope Operator: E. Lundin Instrument: LBC Blue Waiting for...
r4 - 21 Jul 2009 - 22:55 by NormCushing
13-14Feb2007
PartnerObserving Log for 2007 Feb 12/13 Observer: Tom Herbst, Giacomo Beccari Inst. Support: John Hill TOs: Ed Lundin David Gonzalez Fog blowing over the summit...
r6 - 21 Jul 2009 - 22:55 by NormCushing
13June2007
PartnerObserving Log for 2007 June 13/14 Observers: V. Burwitz, D. Thompson Telescope Operator: A. Ceranski Instrument: LBC Blue 23:50 initialized camera! same...
r6 - 21 Jul 2009 - 22:55 by NormCushing
13Mar2007
PartnerObserving Log for 2007 Mar 13/14 Observers: Rick Pogge, Peter Garnavich Inst. Support: John Hill TO: Ed Lundin UTC Mar 13 2100: LBC shutter is still unrepo...
r6 - 21 Jul 2009 - 22:55 by NormCushing
14-15Feb2007
PartnerObserving Log for 2007 Feb 13/14 Observer: Tom Herbst, Giacomo Beccari Inst. Support: John Hill TOs: Ed Lundin David Gonzalez Serious winter storm all af...
r7 - 21 Jul 2009 - 22:55 by NormCushing
14Feb2007
PartnerObserving Log for 2007 Feb 14/15 Observer: Tom Herbst, Giacomo Beccari Inst. Support: John Hill TOs: Ed Lundin David Gonzalez Winter storm continues...Ac...
r6 - 21 Jul 2009 - 22:55 by NormCushing
14Jun2007
PartnerObserving Log for 2007 June 14/15 Observers: V. Burwitz, D. Thompson Telescope Operator: A. Ceranski Instrument: LBC Blue Weather evening: Very thin ...
r8 - 21 Jul 2009 - 22:55 by NormCushing
14Mar2007
PartnerObserving Log for 2007 Mar 14/15 Observers: Rick Pogge, Peter Garnavich Inst. Support: John Hill TO: Ed Lundin UTC Mar 14 1745: Clock board from the CCD co...
r9 - 21 Jul 2009 - 22:55 by NormCushing
15Feb2007
PartnerObserving Log for 2007 Feb 15/16 Observer: Tom Herbst, Giacomo Beccari Inst. Support: John Hill, Dave Thompson TOs: Ed Lundin David Gonzalez Beautifully ...
r8 - 21 Jul 2009 - 22:55 by NormCushing
15Jun2007
PartnerObserving Log for 2007 June 15/16 Observers: V. Burwitz, J. Hill Telescope Operator: A. Ceranski Instrument: LBC Blue Weather evening: Thin high clou...
r8 - 21 Jul 2009 - 22:55 by NormCushing
15Mar2007
PartnerObserving Log for 2007 Mar 15/16 Observers: Rick Pogge, Peter Garnavich Inst. Support: John Hill TO: Ed Lundin UTC Mar 15 2200: Day crew is replacing the s...
r10 - 21 Jul 2009 - 22:55 by NormCushing
15Oct2007
PartnerObserving Log for 2007 Oct 15 16 Observers: D. Sand and M. Coleman Telescope Operator: D. Gonzalez Huerta Beautiful Clear evening. Not a cloud in the sky....
r7 - 21 Jul 2009 - 22:55 by NormCushing
LinksToUsefulInformation
USEFUL LINKS FOR OBSERVERS LBC Instrument: * LBC Team Instrument Page * LBC ETC * LBC GUI for generating PartnerObserving Blocks SDT PartnerObse...
r5 - 21 Jul 2009 - 22:55 by NormCushing
Log20Jan2007
PartnerObserving Log for 2007 Jan 19/20 snowing whole day Main.fan 20 Jan 2007 Observers: Rick Pogge Xiaohui Fan Inst. Support Astronomer: Dave Thompson TOs...
r12 - 21 Jul 2009 - 22:55 by NormCushing
Log21Jan2007
PartnerObserving Log for 2007 Jan 20/21 Observers: Rick Pogge Xiaohui Fan Inst. Support: Dave Thompson TOs: Aaron Ceranski David Gonzalez 1800 UTC Jan 20: Abo...
r11 - 21 Jul 2009 - 22:55 by NormCushing
Log22Jan2007
PartnerObserving Log for 2007 Jan 21/22 Observer: Rick Pogge Inst. Support: none TOs: Aaron Ceranski David Gonzalez Winter storm started around noon. Wind, sno...
r8 - 21 Jul 2009 - 22:55 by NormCushing
Log23Jan2007
PartnerObserving Log for 2007 Jan 22/23 Observers: Rick Pogge Inst. Support: None TOs: Aaron Ceranski A winter storm dumped about a foot of fresh powdery snow on...
r13 - 21 Jul 2009 - 22:55 by NormCushing
Log24Jan2007
PartnerObserving Log for 2007 Jan 23/24 Observers: Rick Pogge Inst. Support: Olga Kuhn TOs: Aaron Ceranski John Hill is also present tonight. UTC Jan 23 2100: At...
r10 - 21 Jul 2009 - 22:55 by NormCushing
LogisticalInformation
Logistical Information: LBT PartnerObserving Room Phone Numbers * 1 520 626 1466 (Tucson Exchange) * 1 928 428 4286 (Safford Exchange) * 1 928 428 9768 (...
r4 - 21 Jul 2009 - 22:55 by NormCushing
NightLog
Nightly PartnerObserving Logs (UT dates) Feb run * Feb 7, 2008 * Feb 8, 2008 * Feb 9, 2008 * Feb 10, 2008 * Feb 11, 2008 * Feb 12, ...
r28 - 21 Jul 2009 - 22:55 by NormCushing
NightlyPlan
Nightly PartnerObserving Plans (UT dates) Feb run * Feb 7, 2008 * Feb 8, 2008, same as yesterday * Feb 9, 2008 Apr run * Apr 2008 General plan...
r10 - 21 Jul 2009 - 22:55 by NormCushing
ObserverCheckList
SDT PartnerObserving Checklist This page is a checklist of basic tasks from afternoon through morning for a night of SDT observing. Afternoon (usually before 4:3...
r3 - 21 Jul 2009 - 22:55 by NormCushing
ObserverChecklist
Observer's Checklist/Information: Basic Safety Information (read all of this!) * WelcomeInfo * SafetyRules High Altitude Considerations * ConsentForm ...
r7 - 21 Jul 2009 - 22:55 by NormCushing
ObservingLogFor2007Jun11-12
PartnerObserving Log for 2007 June 11/12 Observers: V. Burwitz, D. Thompson Telescope Operator: D. Gonzalez Instrument: LBC Blue UT 00:33 so far bad weather thi...
r5 - 21 Jul 2009 - 22:55 by NormCushing
08Oct2007
PartnerObserving Log for 2007 Oct 08 09 Observers: J. Hill, V. Testa Telescope Operator: A. Ceranski Telescope and Instrument Support (alseep on site) 051 25 ...
r6 - 21 Jul 2009 - 22:55 by NormCushing
09Jun2007
PartnerObserving Log for 2007 June 09/10 Technical Observer: J. Hill Telescope Operator: D. Gonzalez Instrument: LBC Blue Technical observing with a bit of SDT M8...
r8 - 21 Jul 2009 - 22:55 by NormCushing
09May2007
PartnerObserving Log for 2007 May 09/10 Technical Observers: J. Hill, E. Diolaiti, A. DiPaola, R. Speziali Telescope Operator: E. Lundin Technical Support: J. Lit...
r6 - 21 Jul 2009 - 22:55 by NormCushing
09Oct2007
PartnerObserving Log for 2007 Oct 09 10 Observers: V. Testa, J. Hill, Stephane Herbert Fort Telescope Operator: David Gonzalez Huerta Telescope and Instrument S...
r4 - 21 Jul 2009 - 22:55 by NormCushing
10Apr2007
PartnerObserving Log for 2007 Apr 10/11 Observers: John Hill, Mark Wagner Telescope Operator: David Gonzales Instrument: LBC Blue Some more SDT related observin...
r4 - 21 Jul 2009 - 22:55 by NormCushing
10Jun2007
PartnerObserving Log for 2007 June 10/11 Observers: V. Burwitz, J. Hill Telescope Operator: D. Gonzalez Instrument: LBC Blue Open after sunset. T=10.9degC, D= 3....
r6 - 21 Jul 2009 - 22:55 by NormCushing
10Mar2007
PartnerObserving Log for 2007 Mar 10/11 Observer: John Hill Telescope Operator: Ed Lundin Some more SDT related observing during engineering time. Note that I'...
r7 - 21 Jul 2009 - 22:55 by NormCushing
10May2007
PartnerObserving Log for 2007 May 10/11 Technical Observers: J. Hill, A. DiPaola, R. Speziali Telescope Operator: E. Lundin Instrument: LBC Blue Some more SDT rel...
r7 - 21 Jul 2009 - 22:55 by NormCushing
11-12February2007
PartnerObserving Log for 2007 Feb 11/12 Observer: Tom Herbst, Giacomo Beccari Inst. Support: John Hill TOs: Ed Lundin David Gonzalez Winter storm started around...
r3 - 21 Jul 2009 - 22:55 by NormCushing
11Mar2007
PartnerObserving Log for 2007 Mar 11/12 Observer: John Hill TO: Ed Lundin Some more bonus SDT observations in engineering time. As with last night, I'm only usin...
r4 - 21 Jul 2009 - 22:55 by NormCushing
11May2007
PartnerObserving Log for 2007 May 11/12 Technical Observers: J. Hill, A. DiPaola, R. Speziali Telescope Operator: E. Lundin Instrument: LBC Blue After waiting for...
r5 - 21 Jul 2009 - 22:55 by NormCushing
12Jun2007
PartnerObserving Log for 2007 June 12/13 Observers: V. Burwitz, D. Thompson Telescope Operator: A. Ceranski Instrument: LBC Blue 02:38 Sunset 02:47 Dome Opened...
r7 - 21 Jul 2009 - 22:55 by NormCushing
25Feb2007
PartnerObserving Log for 2007 Feb 25/26 Observer: John Hill, Emiliano Diolaiti Inst. Support: John Hill TO: Ed Lundin Mixed in with some engineering activities,...
r2 - 21 Jul 2009 - 22:55 by NormCushing
26Feb2007
PartnerObserving Log for 2007 Feb 26/27 Observer: John Hill, Emiliano Diolaiti Inst. Support: John Hill TO: Ed Lundin Mixed in with some engineering activities, ...
r3 - 21 Jul 2009 - 22:55 by NormCushing
ApplyForTime
LBT PartnerObserving Time Proposals Proposals for observing time are currently not being accepted. Proposal Submission * Next proposal due date: * Use thes...
r3 - 21 Jul 2009 - 22:55 by NormCushing
ArizonaQueue
Programs and Progress: * Jan 2009 Run Plan * May Run Summary * April Run Summary * Feb Run Summary * PartnerObserving Schedule * Program Titles ...
r12 - 21 Jul 2009 - 22:55 by NormCushing
December20UT
20 December 2008 UT Summary The humidity dropped and skies cleared but there was too much ice on the enclosure to open. In any case, there was much closed dome t...
r7 - 21 Jul 2009 - 22:55 by NormCushing
December21UT
21 December 2008 UT PartnerObserving Log morning afternoon: * Fernando Vincenzo noticed this morning that the Blue Science controller fans, which they fo...
r6 - 21 Jul 2009 - 22:55 by NormCushing
December22UT
PartnerObserving Log morning Andrea, Fernando Vincenzo wrapping up. In the User Interface, Andrea changed the OB Progress to show number of images taken out of ...
r3 - 21 Jul 2009 - 22:55 by NormCushing
December23UT
23 December UT "PartnerObserving" Log Clouds, humidity, snow and high winds. The winter storm came earlier than expected and there was no chance to open today. ...
r3 - 21 Jul 2009 - 22:55 by NormCushing
December24UT
No PartnerObserving The second winter storm came in last night and there is snow and wind. No chance of opening. Based on the data taken 22 Dec (UT), shims were ...
r3 - 21 Jul 2009 - 22:55 by NormCushing
DitherPatterns
Dither Patterns There are various dither patterns available to select in the LBC GUI: random, circular, rectangular, LBC 4 step and 10 step default and user defin...
r3 - 21 Jul 2009 - 22:55 by NormCushing
FebSummary
0. Summary: Overall a productive run, and certainly way beyond my (modest) expectations. Weather was good; we observed all six nights; had good ( =0.5") seeing o...
r2 - 21 Jul 2009 - 22:55 by NormCushing
ImageGallery
#TopTop Image Gallery This page contains 'typical' images along with some examples of 'odd' images with a brief explanation of what is causing the problem and s...
r13 - 21 Jul 2009 - 22:55 by NormCushing
InstrumentIssues
Issues that LBC Observers Should Know About Last update: 7 February 2008 Data Analysis 1. Prior to the February 2007 SDT time, the exposure times in the head...
r13 - 21 Jul 2009 - 22:55 by NormCushing
03Oct2007
PartnerObserving Log for 2007 Oct 03 04 Observers: J. Hill, V. Testa Telescope Operator: A. Ceranski Never opened because of high humidity and eventually fog. T=...
r3 - 21 Jul 2009 - 22:55 by NormCushing
04Oct2007
PartnerObserving Log for 2007 Oct 04 05 Observers: J. Hill, V. Testa Instrument Support: R. Bertram Telescope Operator: A. Ceranski This is scheduled SDT time whi...
r5 - 21 Jul 2009 - 22:55 by NormCushing
05Nov2007
PartnerObserving Log for 2007 Oct 06 07 Observers: J. Hill, A. Rakich Telescope Operator: D. Gonzalez Huerta (all the technical images aren't included in this log...
r2 - 21 Jul 2009 - 22:55 by NormCushing
05Oct2007
PartnerObserving Log for 2007 Oct 05 06 Observers: J. Hill, V. Testa Instrument Support: R. Bertram Telescope Operator: A. Ceranski First images in 2007106 are th...
r4 - 21 Jul 2009 - 22:55 by NormCushing
06Oct2007
PartnerObserving Log for 2007 Oct 06 07 Observers: J. Hill, V. Testa Instrument Support: R. Bertram Telescope Operator: A. Ceranski 005954 010215 5 bias frame...
r7 - 21 Jul 2009 - 22:55 by NormCushing
07Oct2007
PartnerObserving Log for 2007 Oct 07 08 Observers: J. Hill, V. Testa Instrument Support: R. Bertram Telescope Operator: A. Ceranski 00:49 25 bias frames 01:05 ...
r7 - 21 Jul 2009 - 22:55 by NormCushing
08Jun2007
PartnerObserving Log for 2007 June 08/09 Technical Observer: J. Hill Telescope Operator: D. Gonzalez Instrument: LBC Blue Technical observing with a bit of SDT M8...
r7 - 21 Jul 2009 - 22:55 by NormCushing
2009March20UT
20 March 2009 UT LBC Technical PartnerObserving Observers: O. Kuhn (@LBT), J. Hill (Tucson) Telescope Operator: D. Gonzalez Huerta Thin cirrus over all sky at s...
r6 - 21 Jul 2009 - 22:55 by NormCushing
2009March21UT
21 March 2009 UT LBC Technical PartnerObserving Observers: O. Kuhn (@LBT), J. Hill (Tucson) Telescope Operator: D. Gonzalez Huerta Problems with the NFS disk at...
r4 - 21 Jul 2009 - 22:55 by NormCushing
20Jun2007
PartnerObserving Log for 2007 June 20 21 Observer: D. Thompson LBC Support: A. DiPaola, R. Speziali, R. Bertram Telescope Operator: D. Gonzalez 23:00 AdP RS fou...
r4 - 21 Jul 2009 - 22:55 by NormCushing
20Mar2007
PartnerObserving Log for 2007 Mar 20/21 Observers: Rick Pogge, Eichii Egami Support Scientist: none TO: Aaron Ceranski The Vernal Equinox is at 1707 MST today. ...
r6 - 21 Jul 2009 - 22:55 by NormCushing
21Apr2007
PartnerObserving Log for 2007 Apr 21/22 Observers: A. Grazian, R. Gredel. Night Assistant: J. Hill Telescope Operator: D. Gonzales 17:12 Turn on House Keeping Tu...
r10 - 21 Jul 2009 - 22:55 by NormCushing
21Jun2007
PartnerObserving Log for 2007 June 21 22 Observer: J. Hill LBC Support: A. DiPaola, R. Speziali, R. Bertram Telescope Operator: D. Gonzalez More technical observi...
r5 - 21 Jul 2009 - 22:55 by NormCushing
22Apr2007
PartnerObserving Log for 2007 Apr 22/23 Observers: A. Grazian, R. Gredel. Night Assistant: Telescope Operator: D. Gonzales 18:06 Turn on Blue Channel 18:09 Came...
r6 - 21 Jul 2009 - 22:55 by NormCushing
22Jun2007
PartnerObserving Log for 2007 June 22 23 Observer: J. Hill, A. Rakich LBC Support: A. DiPaola, R. Speziali Telescope Operator: D. Gonzalez Open 02:50 UT without i...
r5 - 21 Jul 2009 - 22:55 by NormCushing
22May2007
PartnerObserving Log for 2007 May 22/23 MST Observers: J. Hill Telescope Operator: A. Ceranski Instrument: LBC Blue Peter is gone. Moon is first quarter. John i...
r2 - 21 Jul 2009 - 22:55 by NormCushing
23Apr2007
PartnerObserving Log for 2007 Apr 23/24 Observers: A. Grazian, R. Gredel. Night Assistant: Telescope Operator: D. Gonzales Snow storm. Then high wind blows away ...
r5 - 21 Jul 2009 - 22:55 by NormCushing
23Jun2007
PartnerObserving Log for 2007 June 23 24 Observer: J. Hill LBC Support: A. DiPaola, R. Speziali Telescope Operator: D. Gonzalez Last night of technical observing ...
r7 - 21 Jul 2009 - 22:55 by NormCushing
2009January20UT
2009 January 20 UT LBC PartnerObserving Log Observers are: Marco Pedani, Roberto Ragazzoni, Fernando Pedichini, John Hill, Mark Wagner, Wayne Wack (LBT) and Olg...
r6 - 21 Jul 2009 - 22:55 by NormCushing
2009January21UT
2009 January 21 UT LBC PartnerObserving Log Main.MarcoPedani 22 Jan 2009 Partly cloudy we start at 01:31 UT on sky Check Collimation Table EL= 84 lbcb.2009...
r3 - 21 Jul 2009 - 22:55 by NormCushing
2009June12UT
Parasitic LBC Red PartnerObserving in Parallel to IRTC LFBG for 12 June 2009 UT LBC Observer: J. Hill (Tucson) IRTC Observers: D. Thompson, M. Pedani (mtn) Telesc...
r3 - 21 Jul 2009 - 22:55 by NormCushing
2009June13UT
Parasitic LBC Red PartnerObserving in Parallel to IRTC LFBG for 13 June 2009 UT LBC Observer: J. Hill (Tucson) IRTC Observers: D. Thompson, M. Pedani (mtn) Telesc...
r3 - 21 Jul 2009 - 22:55 by NormCushing
2009May24UT
02:04 LBCB Trackers failure 02:26 Boot successful 02:00 Around this time, DX was not raising finally did at about 02:30 02:49 Test flats (took data but telescop...
2009May25UT
May 25th UT LBC Tech Night 2 Observers: Pedani, Kuhn OSA: Gonzalez Huerta 02:05 Starting taking 25 Biases. Outside T_air=8.4, T_mir=7.9 Scattered Clouds 05:27...
2009June04UT
2009 June 04 UT Technical Night after LBC maintenance engineering run Observers: O. Kuhn (Tucson), and J. Hill (at beginning) Telescope Operator: A. Ceranski S...
2009June15UT
15 June 2009 UT Tech Night 1 Observers: O. Kuhn (mtn) Operator: W. Wack Instrument Support: J. Morris Weather: some clouds around, not photometric. RH low, ~30% b...
2009June16UT
16 June 2009 UT LBC Tech Night 2 Observers: O. Kuhn (mtn), J. Hill (Tucson, 1st half) Operator: W. Wack Instrument Support: J. Morris Preparations afternoon ...
June04List
Copy of file /Repository/allimagelist_20090604. Filename OBname Filter focusoff lbcb.20090604.043318.fits 1 WT10_352 V BESSEL lbcb.20090604.043755.fits...
2009March24UT
2009 May 24 UT 02:04 LBCB Trackers failure 02:26 Boot successful 02:00 Around this time, DX was not raising finally did at about 02:30 02:49 Test flats (took da...
Log090330
Observers: Fuyon Bian and Peter Milne Operator: Aaron Ceranski Summary Did not open due to problem with building rotation. Details All Times UT Fifth Night of ...
Log090329
Summary Did not open due to high winds. Details Observers: Fuyon Bian and Peter Milne Supporting Astronomer: Marco Pedani Operator: Wayne Wack All Times UT Fou...
Log090328
Summary Weather was nice and clear except last hour before sunrise. Seeing was more stable around 1"some time below 1". We had hard time to focus at the beginning...
Log090327
Summary Note that the pointing was off by ~5 arcmin, and we discovered this at 2am while tuning things up for the Yan program. Seeing was unstable the whole night...
Log090326
Summary We were closed most of the night due to high winds and some snow. The right shutter was never able to open, and so we observed with the blue side only. ...
SXandDXModels
20090320 LBC Blue SX and Red DX Collimation plots. Temperature was about 4 5 C. Main.OlgaKuhn 21 Mar 2009 * SXCollimation20090320.plot: Fit to 20090320 SX...
FocusResults
Best Focus Results Bfoc1 Bfoc2 center Ffoc1 0.2313 2.44 Ffoc2 0.2162 2.31 = F best focus 0.220 gfoc1 gfoc2 gfoc3 gfoc4 ifoc1 0.1154 (near ve lim) 2.37 ifoc2 ...
ListofFocusData
List of images used to measure filter focus offsets Red on left and Blue on right. I used only chip2 (but images were with all 4 chips) and usually the region at ...
Log090125
Summary Details Observers: Fuyan Bian Supporting Astronomer: Marco Pedani Operator: Wayne Wack All Times UT Fifth Night of the AZ queue. 2230 Sky looks pretty go...
Log090124
Summary The seeing is bad. Most of The temperature of the telescope is unstable, making the focus change a lot and it is very hard to converge. The red channel ...
Log090123
Summary Details Observers: David Sand, Fuyan Bian, Zoltan Bolag Supporting Astronomer: Marco Pedani Operator: Wayne Wack All Times UT Third night of the AZ queue...
Log090121
Summary Details Observers: David Sand Supporting Astronomer: Marco Pedani Operator: Wayne Wack All Times UT First Night of the AZ queue. 2250 Sky looks pretty ba...
Log090122
Summary Details Observers: David Sand, Fuyan Bian, Zoltan Bolag Supporting Astronomer: Marco Pedani Operator: Wayne Wack All Times UT Second night of the AZ queu...
Jan09Plan
Object list IC 348: 03:44 32:09 Barlog Segue 1: 10:07 16:04 EdO Aries: 02:19 20:10 EdO M3: 13:42 27:22 EdO M92: 17:17 43:08 EdO Bootes: 14:31 32:28 Fan...
December19UT
19 December 2008 UT morning Doug Mark brought the new F972N20 filter back to Tucson to get a better scan of it. afternoon reviewing logs from last night. Suspi...
December18UT
18 December 2008 times are MST, UT = MST 7h morning The new F972N20 filter in the ring does not fit into the hole in the filter wheel 2. For clarification, the...
October22UT
Oct 22 UT 2008 Collimation 015740 1294 191 512 32 294 678 0.92 sent 015950 5258 0 0 0 0 0 0.74 sent 020656 5319 ...
October20UT
20 October UT 18:20 IIF problems LBCstart had to be done since LBC code had been recompiled 18:30 problems with sending a RB ob when red is unticked Z11...
September25UT
25 September UT Afternoon * Andrea changed the software to fix the problem of single chip observations leaving the detector in a bad bias state. Now integrat...
AprilSummary
Below is a brief summary of the April Arizona LBC run. We observed with the LBCs on the LBT for six nights from Apr 6 to Apr 11, 2008. The Arizona observing team...
MaySummary
Summary: Overall a productive run; problems with camera and telescope jump sometimes frustrating; but weather was beautiful and a fair amount of d...
Log080603
01:30 some thin clouds on the horizon. night might not be photometric. 02:18 attempted to move telescope to zenith, but encountered HBS problem. 02:28 moving to S...
Log080602
00:25 telescope at zenith. We were able to turn the camera off and on to revive LBC red. taking test bias frames. looks normal, no pattern visible in red chip!! 0...
Log080601
01:14 25Bias_Bino: lbcb.20080601.011506.fits lbcb.20080601.012831.fits lbcr.20080601.011531.fits lbcr.20080601.011627.fits lbcr froze after 6 frames due to...
Log080531
Summary The observation is smooth tonight. The sky was clear after the observation began. Some of trailed images (due to the guide star is galaxy or telescope jum...
Plan0805
sky calendar Sunset: 1921, twiend: 2102/LST1322, midnightLST: 1620, twibegin: 0333/LST1955, sunrise 0512; note this 18 deg twilight, we can start/end at least 15 ...
Log080530
Summary The observation was more smooth than yesterday, almost no time wasted tonight. Seeing is good during the first half night (0.7 0.9 for red channel), and b...
Log080529
Summary Several camera problems (rotator, temperature ...)were meet. Two hours observation time lost. The seeing is 0.8 1.0" for red channel and 1 1.6" for the ...
ArimassTablesMay08
May 08 Fields * Bootes * Extended Groth Strip * Herc Main.XiaohuiFan 27 May 2008
CalibrationPlanAZ
Basics We need four kinds of calibration data for this run. * biases (in the afternoon) * twilight flat fields (both evening and morning twilights): in filt...
Log080408
Summary Overall a productive night. Seeing 0.7" 1.3". Observed one OB in UMaII, one OB in Herc, six OBs in Bootes and three OBs in EGS. Lost the last hour of the ...
Log080409
Summary High wind; almost no useful data taken Details Observers: Linhua Jiang, David Sand Supporting Astronomer: Marco Pedani Operator: David Gonzalez Huerta Al...
Log080410
Summary Not open due to wind and humidty Details Observers: Linhua Jiang, David Sand Supporting Astronomer: Marco Pedani Operator: David Gonzalez Huerta All Time...
Log080411
Summary High wind, mediocre seeing. Only tried three OBs, the rest of time closed due to wind. Details Observers: Linhua Jiang, David Sand Supporting Astronomer:...
Plan080409-11
For the rest of this run, our fields are: beginning of the night (1) If Sand feels needs to re do some UMa OBs; (2) otherwise, we will start on Q1137 project, whi...
Log080407
Summary Another smooth night. Almost no time lost to telescope/camera. Lost two hours due to high wind. Sky remained to be photometric. Seeing is around 1.0" in t...
Plan080408
sky calendar Sunset: 1845, twiend: 2015/LST0915, midnightLST: 1300, twibegin: 0430/LST1730, sunrise 0550 observers X. Fan, L. Jiang, B. Weiner forecast mostly cle...
Plan080407
sky calendar Sunset: 1845, twiend: 2015/LST0915, midnightLST: 1300, twibegin: 0430/LST1730, sunrise 0550 observers X. Fan, B. Weiner forecast calls for clear skie...
Log080406
Summary 080406 Overall a smooth and productive night. After an initial rebooting around sunset, LBC didn't fail the entire night. There were several small telesco...
Plan080406
sky calendar Sunset: 1845, twiend: 2015/LST0915, midnightLST: 1300, twibegin: 0430/LST1730, sunrise 0550 observers X. Fan, B. Weiner forecast calls for clear skie...
Plan0804
sky calendar Sunset: 1845, twiend: 2015/LST0915, midnightLST: 1300, twibegin: 0430/LST1730, sunrise 0550 for clear night (1) 2015 2245 UMaII; (2) Bootes 2245 00...
AirmassTablesApr08
Apr 08 Fields * Bootes * Extended Groth Strip * Herc * NGC6503 * QSO1137 * SN_2007af * SN_2007ax * UMaII Main.XiaohuiFan 02 Apr 2008
ObservingSchedule
Run 1: Feb 6 11 Fan, Jiang, Feb 5 8; Herbert Ford, Sand ,Feb 8 11) Run 2: Apr 5 10 Fan, Weiner, Apr 5 7; Jiang, Apr 7 10; Sand, Apr 8 10 Run 3: M...
Team
Main.RolandGredel 03 Mar 2008
Log080211
Summary: Band of high clouds at beginning of the night. Details Observers: Dave Sand, St\xE9phane Herbert Fort, Eric Nielson Supporting Astronomer: None :( Ope...
Log080209
Summary: Another productive night. We struggled early again, due the recovery from a power outage. Otherwise, telescope and camera did well. We did spend fair amo...
Log080210
Summary: There were light clouds at the beginning of the night, so no standard stars were taken. Overall a pretty good night, with long stretches with few proble...
OBCompletion
UMaII 080208: UMaII_pos0_Br, UMaI_pos1_Br, UMIaII_pos2_Br 080209: UMaII_pos3_Br,UMaII_pos4_Br,UMaII_pos5_Br,UMaII_pos6_Br,UMaII_pos7_Br 080210: UMaII_pos8_Br, UM...
Plan080209
Observers Xiaohui Fan, Linhua Jiang, Dave Sand, Stephane Herbert Fort Calendar Sunset: 17:58, twiend: 19:21 (LST0417), twiend: 0545 (LST1442) Afternoon Bias...
Log080208
Summary: Overall a productive night. Photometric with OK seeing (0.9 1.5"). Part of first half lost to telescope and LBC troubles; the last 1.5 hours lost to wind...
Log080207
Summary Bad seeing, lot of instrument troubles. Little useful data on SN and GRB. Observers: Xiaohui Fan, Linhua Jiang Support: Marco Pedani, Olga Kuhn Operator D...
Plan080207
Observers Xiaohui Fan, Linhua Jiang Calendar Sunset: 17:58, twiend: 19:21 (LST0412), twiend: 0546 (LST1438) Afternoon BiasTest.ob, Biast25Dark3.ob Twilight ...
ArimassTables
Feb Fields * Bootes * Extended Groth Strip * Herc * NGC3184 * NGC3351 * NGC3368 * NGC4826 * NGC6503 * QSO1137 * SN2006mq * SN2007af * SN2...
ProgramTitle
AZ (Arizona): PI Topic Balog Deep LBC Imaging of Rho Oph Egami Deep LBT...
OBPrepConsiderations
Considerations for OB Preparation Correct Filter List in LBC OB generating GUI The latest release of the OB GUI (21 November 2007, available at http://lbc.oa roma...
ModIBessel
modified I Bessel transmission The plot and text represent only a nominal modified I Bessel filter. * I_Bessell_mod.txt: Nominal modified I Bessel transmissio...
Yfilt
Y filter transmission Note that the Y filter is a PI filter. Please contact Xiaohui Fan (fan #64;as.arizona.edu) for availability. Transmission is measured in a ...
SdssZ
Sloan z' transmission Measured in a collimated beam at 4 corner points (dotted curves) and through the center (solid curve). In the f/1.45 beam, the wavelengths w...
SdssI
Sloan i' transmission Measured in a collimated beam at 4 corner points (dotted curves) and through the center (solid curve). In the f/1.45 beam, the wavelengths w...
SdssRr
Sloan r' transmission Measured in a collimated beam at 4 corner points (dotted curves) and through the center (solid curve). In the f/1.45 beam, the wavelengths w...
17Oct2007
Main.JohnHill 18 Oct 2007 Sand and Coleman here. Beautiful sunset. Candidate Photometric night. Good way to start the last night of SDT EVER. 0110 Taking 5 b...
14Oct2007
Main.JohnHill 15 Oct 2007 Hi Matt Coleman and Dave Sand are at the LBT. Olga is helping and it is great. 0110 UT Opening Dome. Having some trouble retracting....
13Oct2007
Observers: C.S. Kochanek, V. Testa, D. Gonzales Huerta Clouds all around. Humidity 80% at sunset (00:56) 00:54 5bias 005405 005625 01:00 Slew to 1hr E, 2...
12Oct2007
Observers: C.S. Kochanek, V. Testa, D. Gonzales HuertaClear sky but with a few thin clouds at horizon toward East. Opened at 01:00 01:10 Azimut encoders can't b...
11Oct2007
Observers: V. Testa, S. Herbert Fort, D. Gonzales Huerta Quite nice and sunny in the afternoon. Sky looks very clear.At the moment it is candidate photometric. 06...
10Oct2007
Observers are: V. Testa, S. Herbert Fort, D. Gonzales Huerta Mostly sunny with few thin clouds in the afternoon 00:35 25 bias 003632 005038 01:00 Sunset Stil...
16Sep2007
19:45 Stop and restart firefox on lbtmu03, since it was taking up to 62% memory and was slow to respond. In new firefox, start LBC and run up. 19:50 Start series ...
UserDitherPatterns
The LBC focal plane consists of four edge butted CCDs with gaps of ~18 arcseconds between the arrays. You can find more details about LBC on its instrument page:...
17May2007
17May2007 Observer: Peter Buschkamp, Nicolas Bouche TO: David Gonzalez 20:00 we are ready, LBC is ready, the telescope seems ready, but rain is coming our way 21:...
ObserverSchedule
Current SDT Observer Schedule Date Observer #1 Observer #2 LBTO Science LBTO Technical LBTO Software June 10 Vadim Burvitz (MPE) ??...
21May2007
Last night of the May SDT run. In short: Lots of moonlight and bad seeing after a while. High wind speed with gusts 16m/s. Humidity too high after 2am. closed ...
20May2007
Observer: Peter Buschkamp Telescope Operator: David Gonzalez What got observed: 1. Flats in U spec (0 and 180deg) 2. Flats in R sloan (0 and 180deg) 3. Flats in B...
19May2007
20:09 last threatening cloud has moved away and we are opening. 21:05 standard ; sky clear; seeing 0.8" 21:25 going to GRB 070419 seeing 0.8"; sky clear 22:25 M81...
18May2007
20:00 Cloud overhead dropped rain drops. Otherwise, horizon clearing up. 20:40 we are open! Not photometric. 20:55 LBC archive stuck. No data being put in /Repo...
MayStandards
An updated set of standards for May is at /home/sdtobs/1SDTOBs/Standards_May. Following is an airmass table for all of these fields: field air eve twi...
InstrumentHomePage
LBTO Instrumentation Home Page General Acknowledgements: The LBTO requests that publications based on data obtained at the LBT include the following statement as...
11Apr2007
Observers: Jason Harris, Paul Martini, Jose Prieto TO: David Gonzalez Instrument Support: John Hill Almanac: Sunset ( 3000 m horizon): 18 55 MST; Sunrise: 5 45 ...
15Apr2007
$ 18:50: Welcome to LBT, Dr. Sand! High winds, but it should be okay. $ 19:10: Twilight flats Filter scale counts U 1.0 30000, 25000, * V ...
r3 - 16 Apr 2007 - 09:37 by JasonHarris
14Apr2007
$ 16:15: 25 Bias frames and 900 sec Dark $ 18:45: Sky is mostly clear. Some wispy clouds in the northwest, so it's not photometric. $ 19:00: Doing Twilig...
r6 - 15 Apr 2007 - 12:20 by JasonHarris
ObsTools
Planning Tools * airmass vs time plot tool from ING LBT coordinates: E longitude deg and latitude deg and altitude m are 250.111 32.7013 3221 Note that this...
13Apr2007
$ 16:45: Biases and 900 sec Dark $ 18:30: The fog has lifted, but it's still mostly couldy, and humidity is 98%. The satellite images look promising, though. ...
r6 - 14 Apr 2007 - 11:02 by JasonHarris
ObTracking
Tracking OB execution During the night, one of the SDT observers should be monitoring the incoming data, checking that the image quality is good and the seeing ...
MarchStandards
Standard Star Fields for March 2007 Prepared by Ed Olszewski, David Sand, and Jill Bechtold 2007 March 12 This is the list of photometric standard fields for th...
CalibrationPlan
CALIBRATION OBSERVATIONS DURING SDT I: Background By: Edward Olszewski, David Sand, Jill Bechtold, Dave TrillingLast update: Jan. 12, 2007 Summary: This p...
ObAssociation
Mapping from queue listing to OB name Under construction... some notes follow. Mapping from 'queue scheduler' listing to name of OB in /home/sdtobs/1SDTOBs/ qu...
24Feb2007
PaulSmith? 24 Feb 2007 SDT? Observer1: Nicolas Bouche, Peter Buschkamp Observer2: Paul Smith Inst. Support: TOs: Ed Lundin and David Gonzalez 1730 1800: biases ...
23Feb2007
Main.PaulSmith 23 Feb 2007 SDT Observer1: Nicolas Bouche, Peter Buschkamp Observer2: Paul Smith Inst. Support: TO: Ed Lundin Blizzard during the afternoon and ea...
20Feb2007
Main.PaulSmith 20 Feb 2007 SDT Observer1: Nicolas Bouche, Peter Buschkamp Observer2: Paul Smith Inst. Support: Olga Kuhn TO: Aaron Ceranski 1820: Start biases. ...
21Feb2007
Main.PaulSmith 21 Feb 2007 SDT Observer1: Nicolas Bouche, Peter Buschkamp Observer2: Paul Smith Inst. Support: Olga Kuhn TO: Aaron Ceranski Some cirrus to W and ...
22Feb2007
Main.PaulSmith 22 Feb 2007 SDT Observer1: Nicolas Bouche, Peter Buschkamp Observer2: Paul Smith Inst. Support: Olga Kuhn TO: Ed Lundin 1835: Flats UBVgr (0 deg PA...
17Feb2007
Observer1, Nicolas Bouche, Peter Buschkamp Observer2, Giacomo Beccari Inst. Support, Dave Thompson TOs; David Gonzalez, Aaron Ceranski Nice sunset (means lots of ...
ConnectingToPrinters
Printing at the LBT from your Laptops. There is a standard 600dpi black and white HP Laser Jet printer in the control room for use by visiting astronomers. Once ...
SDTDataTransfer
SDT Data Transfer to Tucson for Dissemination: In the spirit of the SDT observations, the observing team members are requested to NOT copy the data to their lap...
SafetyRules
Basic Safety Rules Elevation Altitude Whether you are an experienced trekker, a visitor or an observer, altitude does not discriminate. Altitude effects the you...
r5 - 08 Feb 2007 - 21:17 by GriselaKoeppen
LUCIFERHomePage
#TopTop LUCIFER: LBT Utility Camera ( Spectrograph) Imaging Fields for Extragalactic Research. Overview Quick links to: Main.DavidThompson 07 Feb 2007
InstrumentSupportGroup
LBTO Instrumentation Support Group Personnel Contact any of the following persons for questions or problems with LBT instrumentation or concerning science observa...
RoadConditions
Information on Mountain Road Conditions. Recently, two LBTO staff members got into difficulties driving down the snowy mountain road at night. They were unable...
Log25Jan
Main.okuhn 27 Jan 2007 * 20070125.hselect: 24 25 Jan 2007 Image Listing
NEW - 27 Jan 2007 - 01:04 by okuhn
Log24Jan
hselect /Repository/20070124/lbcb.*.fits 1 $I,lbcobname,objname,imagetyp,filter,exptime,telra,teldec 20070124.hselect Image_name OBname Imagetyp FILTER ExpTime...
NEW - 27 Jan 2007 - 01:02 by okuhn
QuickLookTools
Quick Look Data Inspection Tools Once data have been acquired, there are some simple image inspection tasks you can perform to quickly assess the image quality. ...
r4 - 24 Jan 2007 - 02:00 by pogge
KeyRadioDistribution
To insure smooth access to the observatories and to insure your safety during your LBTO visit, all personnel must check in at the Base Camp to: * Log their pre...
IRAF4dummies
Dealing with LBC data using IRAF and ds9 (for Dummies) You must have a current version of ds9 and IRAF. At the moment, at least at Steward, not everybody ha...
r6 - 20 Jan 2007 - 22:34 by pogge
Log19Jan2007
Main.jill 16 Jan 2007 Main.gredel 19 Jan 2007 covered all day, getting a bit better towards sunset; low wind 3 4 m/s; humidity 50 60% Ambient temperature ...
r3 - 19 Jan 2007 - 04:00 by gredel
Log18Jan2007
Main.jill 16 Jan 2007 Main.gredel 18 Jan 2007 cirrus, high clouds moving in, low humidity 18%, how wind pointer has been changed again to /Repository, let's s...
MeasuredSkyinV
The only data taken so far where sky is easily measured are the Ryan exposures on 15Jan2007, where the exposure time is 240 sec andthe V filter was used. Note tha...
NEW - 18 Jan 2007 - 00:02 by jill
Log17Jan2007
Main.jill 15 Jan 2007 Main.gredel 17 Jan 2007 work on rotator all day, possible cause identified, command rotator spinning motor shaft found, but gear not...
r5 - 17 Jan 2007 - 10:32 by gredel
Log16Jan2007
Main.jill 13 Jan 2007 Main.gredel 15 Jan 2007 afternoon, no wind no clouds no ice sorry, lots of ice on the roof. Snow melting heaters not very efficient, ...
r4 - 16 Jan 2007 - 17:47 by gredel
Log15Jan2007
Main.jill 13 Jan 2007 Main.gredel 14 Jan 2007 clear sky, low wind, strong ice accumulation on enclosure, trying to de ice by exposing walls to sun Mirror at ...
r3 - 15 Jan 2007 - 01:38 by gredel
TwilightObservations
During Evening and Morning Twilight: SUMMARY: 1. Evening Twilight: Take twilight flats, then let the telescope...
r6 - 14 Jan 2007 - 23:23 by jill
Log14Jan2007
Main.jill 13 Jan 2007 Main.gredel 13 Jan 2007 20:00 in the fog, enclosure covered with ice (except of roof, de iced) 22:00 fibre replacement wraps etc. 22:...
r3 - 14 Jan 2007 - 23:14 by jill
ObserverJobs
Here is one of numerous jobs that we expect LBT observers to do on the mountain. Fortunately, Roland Gredel the first SDT observer is experienced at both obs...
Log13Jan2007
Main.gredel 12 Jan 2007 22:15 high wind strong gusts, clouds 22:24 25x bias 800 ADU 13 Jan 00:00 UT strong wind gusts, fog, enclosure covered with ice 03:...
NEW - 13 Jan 2007 - 16:51 by jill
AfternoonActivities
UNDER CONSTRUCTION The LBTO will be carrying out engineering until 4:30 pm every day. At 4:30 pm, have the staff darken the dome, and carry out the following ...
r2 - 12 Jan 2007 - 20:16 by jill
L95info
Photometric/Astrometric Standard Field L95 J2000 RA,Dec = 03:53:11 00:00:00 Exposure Times:1 sec and 2 sec in BVgr 1, 4, 10 sec in U 44 U stan...
NEW - 11 Jan 2007 - 22:18 by jill
MountainWorkstations
Mountain Workstations Currently the control room has five workstations available, one of which is primarily used by the telescope operator (on the far right as yo...
ProgramTitles
Project lead is responsible for OB preparation. Hard copies of the proposals are available on the mountain, and in Tucson. These are not for public distributio...
r2 - 11 Jan 2007 - 00:55 by jill
PemitsInformation
MT. GRAHAM RED SQUIRREL BRIEFING: The endangered Mt. Graham red squirrel is protected by the Federal Endangered Species Act of 1973, as amended. It is illegal to ...
r2 - 10 Jan 2007 - 23:51 by GriselaKoeppen
EmergencyInformation
Emergency Information In Case of Emergency on the Mountain: * 24 hour Emergency Number: 911 * Univ of Arizona Police Department Dispatch: 1 520 621 8273 ...
CreateNewSDTTopic
Main.DavidThompson 26 Dec 2006
r2 - 09 Jan 2007 - 17:48 by jill
DrivingToTheLBT
Driving to the LBT. LBTO is located on top of Mt. Graham, Arizona, approximately 150 highway miles from The University of Arizona in Tucson (see map).link here. ...
PhoneUsage
Phone Usage at the LBTO Cell phone coverage at the observatory is spotty at best, do not count on that as a reliable phone connection. LBTO Phone list (both of ...
LodgingReservationsSafford
Hotels in the Safford area: Best Western Days Inn 1391 West Thatcher Blvd. 520 East Hwy 70 (928) 428 0521 (928) 428 5000 Comfort Inn Quality I...
NEW - 05 Jan 2007 - 22:34 by GriselaKoeppen
MealsattheObservatory
Meals at the Observatory Food is always an important topic and here we have good and bad news. The good news is that you may have any food that you would like wh...
r2 - 05 Jan 2007 - 22:27 by GriselaKoeppen
LodgingReservationsTucson
Lodging Reservations in Tucson. Hotels in the Tucson area: ADOBE ROSE INN 940 N. Olsen Ave. Tucson, AZ 85719 1 800 328 4122 / (520) 318 4644 SHERATON TUCSON HO...
r2 - 05 Jan 2007 - 22:06 by GriselaKoeppen
ArrivingatTucsonAirport
Arriving at Tucson International Airport You will arrive at Tucson International Airport. You must have a valid passport and a U.S. Visa in order to be allowed t...
WelcomeInfo
Welcome Message from our Program Manager This information should be distributed where appropriate We are very happy that you will soon be coming to the Large Bino...
SubjectXforSDT
SDT Data Transfer to Tucson for Dissemination: In the spirit of the SDT observations, the observing team members are requested to NOT copy the data to their lapto...
ConnectingLaptops
Connecting Laptops to the LBT Astronomer's Net. The LBT operates under the paradigm that every Astronomer visiting the observatory will arrive with a laptop compu...
LbcIrafPackage
LBT IRAF Packages Some tasks which are useful to list, display and run a basic reduction while at the telescope have been incorporated into the packages LBTtools...
NEW - 23 Dec 2006 - 09:29 by okuhn
ConsentForm
LBTO Observer's Consent Form Add content here Main.DavidThompson 22 Dec 2006
TransmissionUBVgR
* Transmission curves for UBVg'r': Solid curves are nominal transmission curves. For the Sloan filters these correctly show the blueshift that results when these ...
NEW - 21 Dec 2006 - 09:00 by okuhn
TransUBVgR
* Transmission curves for UBVg'r': Main.okuhn 21 Dec 2006
NEW - 21 Dec 2006 - 08:56 by okuhn
VbesselA
V Bessel transmission Nominal transmission data and plot; not measured for actual filter in LBC Blue. * LBCB_Bessel_V.txt: V Bessel transmission table: lambda ...
NEW - 21 Dec 2006 - 08:29 by okuhn
BbesselA
B Bessel transmission Nominal transmission data and plot; not measured for actual filter in LBC Blue. * LBCB_Bessel_B.txt: B Bessel transmission table: lambda ...
NEW - 21 Dec 2006 - 08:29 by okuhn
UbesselA
U Bessel transmission Nominal transmission data and plot; not measured for actual filter in LBC Blue. * LBCB_Bessel_U.txt: U Bessel transmission table: lambda ...
NEW - 21 Dec 2006 - 08:28 by okuhn
SdssG
Sloan g' transmission There are three tables: one is for the nominal filter, computed assuming a convergent f/1.45 beam as in LBC Blue and the other two (labelled...
NEW - 21 Dec 2006 - 08:26 by okuhn
SdssR
Sloan r' transmission There are three tables: one is for the nominal filter, computed assuming a convergent f/1.45 beam as in LBC Blue and the other two (labelled...
NEW - 21 Dec 2006 - 08:25 by okuhn
LbtAlmanac
LBT Almanac * LBT2007.pdf: LBT2007.pdf * LBT2006.pdf: LBT2006.pdf LBT coordinates On the LBT (mountain) machines skycalc is hardwired (not yet) to automati...
NEW - 18 Dec 2006 - 22:37 by okuhn
WeatherLinks
Weather Information Mt. Graham and the LBT * Live view of the LBT * NWS Forecast for Mt. Graham * Infrared Satellite Loop * Local Radar Loop ...
r2 - 03 Nov 2006 - 19:49 by duplessis
PlanningYourRun
Main.AlexLovellTroy 01 Nov 2006
r2 - 02 Nov 2006 - 00:13 by AlexLovellTroy
ObservationRegistration
Name Type Size Values Tooltip message Attributes Full Name text 30 Observer J. Astronomer your full name e mail address text 30 ...
r2 - 02 Nov 2006 - 00:05 by AlexLovellTroy
LodgingInformation
Testing...1...2...3 Main.DavidThompson 24 Oct 2006
Number of topics: 454
< Previous Page 2 of 2

See also: rss-small RSS feed, recent changes with 50, 100, 200, 500, 1000 topics, all changes
Topic revision: r1 - 26 Feb 2018, ProjectContributor
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback